OpenStudy (anonymous):

Find the exact value of cos^1 -sqrt3/2

OpenStudy (anonymous):

when does cosine take a value of sqrt (3/2) ?

OpenStudy (anonymous):

Idk I'm completely lost & just trying to get an answer

OpenStudy (er.mohd.amir):

since cos x=sqrt3/2 give at x=30 degree but its -ve so ans is 180-30=150

OpenStudy (anonymous):

How'd you come up with that answer